Divide the following. 5,730,000 ÷ 300


1
Expert's answer
2021-07-01T11:48:17-0400

Solution.


"5730000:300=57300:3"

Since the sum of the digits of the number 57300 is equal to 5+7+3=15 and divisible by 3, then the number 57300 is divisible by 3.

57300=30000+27000+300.

From here 57300:3=(30000+27000+300):3=30000:3+27000:3+300:3=

10000+9000+100=19100.

We will have 57300:3=19100.

Answer. 19100.